# ActivationReasoning: Logical Reasoning in Latent Activation Spaces

### v1

@@ Abstract
Large language models (LLMs) excel at generating fluent text, but their internal reasoning remains opaque and difficult to control. Sparse autoencoders (SAEs) make hidden activations more interpretable by exposing latent features that often align with human concepts. Yet, these features are fragile and passive, offering no mechanism for systematic reasoning or model control. To address this, we introduce ActivationReasoning (AR), a framework that embeds explicit logical reasoning into the latent space of LLMs. It proceeds in three stages: (1) Finding latent representations, first, latent concept representations are identified (e.g. via SAEs) and organized into a dictionary; (2) Activating propositions, at inference time, AR detects activating concepts and maps them to logical propositions; and (3) Logical reasoning, applying logical rules over these propositions to infer higher-order structures, compose new concepts, and steer model behavior. We evaluate AR on multi-hop reasoning (PrOntoQA), abstraction and robustness to indirect concept cues (Rail2Country), reasoning over natural and diverse language (ProverQA), and context-sensitive safety (BeaverTails). Across all tasks, AR scales robustly with reasoning complexity, generalizes to abstract and context-sensitive tasks, and transfers across model backbones. These results demonstrate that grounding logical structure in latent activations not only improves transparency but also enables structured reasoning, reliable control, and alignment with desired behaviors, providing a path toward more reliable and auditable AI.
@@ 1 Introduction
Large language models (LLMs) demonstrate remarkable abilities in semantic disambiguation, knowledge retrieval, and generative tasks (Brown et al., 2020; OpenAI, 2023). Yet, their internal activations are distributed and entangled, with multiple unrelated features encoded in overlapping representational dimensions. This phenomenon, known as superposition, obscures conceptual representations and limits both interpretability and intervention Marcus (2020); Bender et al. (2021); Elhage et al. (2022). More fundamentally, autoregressive transformers lack explicit propositional structure, making systematic reasoning, compositional generalization, and rule enforcement difficult (Elhage et al., 2022; Bommasani et al., 2021; Woydt et al., 2025; Delfosse et al., 2025). These challenges extend beyond transparency to robustness, controllability, and safe deployment in real world (Gabriel, 2020; Wei et al., 2022). Mechanistic interpretability, and particularly sparse autoencoders (SAEs), promises remedy by surfacing latent features that often align with human-recognizable, monosemantic concepts (Bricken et al., 2023; Templeton et al., 2024). However, SAEs alone remain limited; features can still be polysemous, contextually unstable, or overly low-level (Leask et al., 2025; Härle et al., 2025; Peng et al., 2025). Critically, SAEs lack mechanisms for composition and higher-order reasoning.
Formal logic, in contrast, offers explicit compositionality, well-defined inference rules, and transparency in how conclusions are derived (Lloyd, 1984; Russell & Norvig, 2009; Shapiro & Kouri Kissel, 2024). Yet, logic-based reasoning presupposes discrete propositional units, whereas LLMs rely on continuous, superposed representations that obscure such structure (Marcus, 2020; Bender et al., 2021; Stammer et al., 2024). While this mismatch has long hindered direct integration, recent progress with SAEs provides a concrete foothold by inducing monosemantic features that approximate discrete propositional units required for symbolic reasoning.
@@
To this end, we propose ActivationReasoning (AR), a framework that embeds explicit logical reasoning into the latent space of LLMs. AR operates in three stages (see Fig.  2): (1) Finding latent representations, where atomic concepts are grounded in SAE-derived features; (2) Activating propositions, which monitors their activations during inference; and (3) Logical reasoning, which composes these propositions into higher-order concepts through user-defined rules. This pipeline extends SAEs from feature extraction to structured reasoning and control. As illustrated in Fig.  1, AR can recover composite concepts absent from the SAE space (e.g., Golden Gate Bridge from Bridge, San Francisco, and US); disambiguate polysemous features such as different uses of red; enforce safety constraints directly in latent space; and perform multi-step reasoning. On the analysis side, this enables fine-grained inspection of model behavior, such as tracing failure cases to specific activations, probing compositional generalization, or evaluating adherence to safety rules. On the control side, AR supports direct interventions during inference by amplifying or suppressing activations to steer the model toward desired behaviors.
Our contributions are twofold: (i) we introduce ActivationReasoning (AR), a framework that embeds logical reasoning into the latent space of LLMs by treating SAE features as propositions and composing them through explicit rules; and (ii) we demonstrate through experiments on reasoning, meta-concept generalization, and safety that latent activations serve as a viable substrate for structured logical reasoning, enhancing transparency, robustness, and control in LLMs.
@@ 3 ActivationReasoning
ActivationReasoning (AR) enables logical reasoning directly within the latent activations of LLMs (Fig.  2), yielding interpretable, controllable, and robust behavior in downstream tasks. The process has three stages: (i) Finding latent representations, where concept representations are identified in the LLM’s latent space and collected in a dictionary; (ii) Activating propositions, where token-wise activations are detected at inference time and treated as propositional units; and (iii) Logical reasoning, where rules are applied over the activated propositions to infer higher-order structures, compose new concepts, and steer model behavior. As a running example, we infer the concept Golden Gate Bridge, which is not initially present in the dictionary but can be derived from the joint activation of three base concepts: Bridge, San Francisco, and USA.
@@ 3.1 Finding Latent Representations
The first stage of AR (cf. Fig.  2, top half) constructs a dictionary $D$ of interpretable concepts, each grounded in the latent space of the LLM. Every concept (e.g., Bridge, USA) is defined by a concept name, latent concept representation, and a soft threshold.
@@
Towards Interpretable Activations. Raw LLM activations are highly entangled and subject to superposition, encoding multiple features in overlapping dimensions (Elhage et al., 2022), which obstructs recovery of monosemantic representations (Park et al., 2023). Logical reasoning, however, requires such representations as propositional building blocks. Thus, we project activations into a sparse and more interpretable feature space using SAEs (Bricken et al., 2023), which promote monosemanticity by aligning individual dimensions with distinct, human-interpretable concepts. The framework remains method-agnostic and can directly integrate future advances in interpretability.
@@
Types of latent representations. Building on SAE features, we propose three forms of latent representation $r\in\{\mathcal{R}_{\text{single}},\mathcal{R}_{\text{multi}},\mathcal{R}_{\text{relation}}\}$ (cf. Fig.  2, top right). (1) A Single-feature representation ($\mathcal{R}_{\text{single}}$) ties a concept to one SAE feature, reflecting the common assumption that some features can align closely with human-interpretable notions (Bricken et al., 2023). (2) A Multi-feature representation ($\mathcal{R}_{\text{multi}}$) instead associates a concept with a set of $k$ SAE features. This relaxes the strict monosemantic assumption, acknowledging that features often capture only fragments of a concept or exhibit polysemy (Leask et al., 2025; Härle et al., 2025). (3) A Relational-feature representation ($\mathcal{R}_{\text{relation}}$) defines a concept by a set of relations among SAE features. In practice, we instantiate this via a shallow decision tree, which balances expressiveness with interpretability. Simple concepts like Bridge or USA can be captured by $\mathcal{R}_{\text{single}}$ or $\mathcal{R}_{\text{multi}}$, while abstract or contextual notions benefit from $\mathcal{R}_{\text{relation}}$. E.g., the broader concept hate may be recognized if features linked to slurs or demeaning stereotypes are activated, while benign uses of similar words are filtered out.
@@
Constructing Representations. Representations can be obtained either manually or automatically. In the manual case, a human expert defines $r_{c}$ by assigning SAE encoder features $(\mathrm{SAE}_{E})$ to concepts. For example, features consistently activating for the notion of a Bridge can be directly assigned.
@@
Concept dictionary output. The output of this stage is a dictionary $D=\{(n_{c},r_{c},\tau_{c})\}_{i=1}^{N}$. Each entry contains a concept’s name, representation, and threshold. For our running example, $D$ includes the base concepts Bridge, San Francisco, and USA, which will later compose into the higher-level concept Golden Gate Bridge.
@@ 3.2 Activating Propositions
For $\mathcal{R}_{\text{single}}$, the score corresponds to the value of a single SAE feature (cf. Fig.  2, purple box). For $\mathcal{R}_{\text{multi}}$, evidence is aggregated from multiple indices using a weighting vector $w$ (e.g., log decay), normalized to sum to one, i.e., forming a convex combination. For $\mathcal{R}_{\text{relation}}$, the score is obtained by traversing the decision tree $r_{c}$ on $l_{t}$, producing a probability of concept presence.
@@ 3.3 Logical Reasoning
Reasoning. Having established a set of user-defined rules, reasoning proceeds via forward chaining. Starting from the initial activation matrix $A$ of logical propositions, inference rules are applied iteratively to derive new conclusions, until a fixed point is reached where no further inferences can be drawn (Russell & Norvig, 2009). We formalize the explicit semantics in App.  D. In practice, rules may operate over $A_{\text{local}}$ and $A_{\text{global}}$. The former provides evidence for explicit concepts such as Bridge, San Francisco, and USA at specific tokens, while the latter captures more abstract notions in the sequence, such as sarcasm. For reasoning, we feed discretized activations $A$ into the rule engine (e.g., $A_{\text{local}}[\textit{Bridge},15]=0.9$ is treated as the Boolean proposition Bridge). Applying the compositional rule above to the evidence allows us to infer the higher-level proposition Golden Gate Bridge, even if no SAE feature directly encodes this concept.
Enriched matrix. We denote the resulting enriched matrix of logical propositions as $A^{\prime}$ (cf. Fig.  2, bottom right). In contrast to the raw activations $A$, which contain only directly activated concepts, $A^{\prime}$ integrates both detected and inferred propositions, providing a structured and interpretable representation that can be leveraged for downstream analysis and control.
@@ 3.4 Integration into Downstream Tasks
For analysis, $A^{\prime}$ allows fine-grained inspection of model behavior, such as tracing failure cases to specific concept activations (cf. Sec.  4, Rail2Country), probing compositional generalization (cf. Fig.  1) or evaluating adherence to safety constraints (cf. Sec.  4, safety experiments).
@@ 4 Experimental Evaluations
Experimental Setup. We evaluate AR on two backbone models, Llama-3.1-8B (AI@Meta, 2024) with EleutherAI’s SAE (EleutherAI, 2024) attached after layer 23, and Gemma-2-9B (Team, 2024) with a Gemma-Scope SAE (Lieberum et al., 2024) attached after layer 20. All experiments use greedy decoding. For evaluation, we apply $\mathcal{R}_{\text{multi}}$ in the first three benchmarks and compare $\mathcal{R}_{\text{single}}$, $\mathcal{R}_{\text{multi}}$, and $\mathcal{R}_{\text{relation}}$ in the fourth. We report results against several baselines: the same backbones without AR, larger instruction-tuned models (Llama-3.1-70B and Gemma-2-27B), commercial models (GPT-4o (OpenAI, 2023)), and the reasoning model DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025). Additionally, we provide comparison to further IT and CoT baselines in App.  E and analysis of runtime efficiency in App.  B. Baseline LLMs generate answers directly, whereas AR performs reasoning in the activation space and steers the downstream generation. Details on datasets and hyperparameters are noted in App.  A.
To study this setting, we introduce Rail2Country (R2C), a benchmark for reasoning over meta-level concept descriptions, inspired by earlier train-based reasoning tasks (Michalski, 1980; Mitchell, 1997; Helff et al., 2024; 2025). Each instance consists of a context that lists countries with their flag color sequences and a query, a short train description whose sequence of car colors must be mapped to the corresponding country (e.g., red–white–blue $\rightarrow$ France). R2C comes with two versions. In R2C-Mono, colors are explicitly stated in the trains’ descriptions, producing clear SAE activations (Bricken et al., 2023; Templeton et al., 2024). In R2C-Meta, explicit color mentions are replaced with similes (e.g., ’colored like a tomato’ for red). In this way, a color cue is entangled with object semantics (’tomato’), contextual phrasing (’like’), and world knowledge (tomatoes are red). Such descriptions are expected to distribute activation across multiple SAE features (Elhage et al., 2022). Further details and examples are provided in Sec.  A.2.1. For evaluation, we initialize AR with $\mathcal{R}_{\text{multi}}$ using the R2C-Mono train set and encode the context rules into $L$.
@@
Towards Real-World Reasoning Scenarios. Next, we extend our evaluation to ProverQA (Qi et al., 2025), a benchmark of LLM-generated problems that present naturalized reasoning tasks which reflect real-world linguistic diversity. It has three complexity tiers, namely, easy, medium, and hard. Each task provides a context of facts and rules as well as a query that is to be proven true, false, or uncertain. For evaluation, we initialize AR with $\mathcal{R}_{\text{multi}}$) using the set of available propositions, and encode the context rules into $L$. Results on ProverQA are summarized in Tab.  1 (right). We observe that baseline LLMs perform poorly. Larger instruction-tuned models and GPT-4o achieve stronger results on the easy tier but deteriorate on medium and hard problems, and the reasoning model DeepSeek-R1-8B follows a similar trend. In contrast, AR yields substantial gains across all tiers, AR (Llama) reaches 93%/91%/71% (easy/medium/hard) and AR (Gemma) 94%/91%/70%, consistently outperforming both the much larger LLMs and reasoning models. On the hard tier, only AR maintains solid performance near 70%, while all other models fall below 50%. This highlights its robustness to linguistically diverse and complex reasoning tasks and demonstrates that structured reasoning in latent space generalizes effectively to natural, increasingly difficult scenarios.

### v2

@@ Abstract
Large language models (LLMs) excel at generating fluent text, but their internal reasoning remains opaque and difficult to control. Sparse autoencoders (SAEs) make hidden activations more interpretable by exposing latent features that often align with human concepts. Yet, these features are fragile and passive, offering no mechanism for systematic reasoning or model control. To address this, we introduce ActivationReasoning (AR), a framework that embeds explicit logical reasoning into the latent space of LLMs. It proceeds in three stages: (1) Finding latent representations, first, latent concept representations are identified (e.g. via SAEs) and organized into a dictionary; (2) Activating propositions, at inference time, AR detects activating concepts and maps them to logical propositions; and (3) Logical reasoning, applying logical rules over these propositions to infer higher-order structures, compose new concepts, and steer model behavior. We evaluate AR on multi-hop reasoning (PrOntoQA), abstraction and robustness to indirect concept cues (Rail2Country), reasoning over natural and diverse language (ProverQA), and context-sensitive safety (BeaverTails). Across all tasks, AR scales robustly with reasoning complexity, generalizes to abstract and context-sensitive tasks, and transfers across model backbones. These results demonstrate that grounding logical structure in latent activations not only improves transparency but also enables structured reasoning, reliable control, and alignment with desired behaviors, providing a path toward more reliable and auditable AI. Code and Dataset available at https://github.com/ml-research/ActivationReasoning
@@ 1 Introduction
Large language models (LLMs) demonstrate remarkable abilities in semantic disambiguation, knowledge retrieval, and generative tasks (OpenAI, 2020; 2023). Yet, their internal activations are distributed and entangled, with multiple unrelated features encoded in overlapping representational dimensions. This phenomenon, known as superposition, obscures conceptual representations and limits both interpretability and intervention Marcus (2020); Bender et al. (2021); Elhage et al. (2022). More fundamentally, autoregressive transformers lack explicit propositional structure, making systematic reasoning, compositional generalization, and rule enforcement difficult (Elhage et al., 2022; Bommasani et al., 2021; Woydt et al., 2025; Delfosse et al., 2025). These challenges extend beyond transparency to robustness, controllability, and safe deployment in real world (Gabriel, 2020; Wei et al., 2022). Mechanistic interpretability, and particularly sparse autoencoders (SAEs), promises remedy by surfacing latent features that often align with human-recognizable, monosemantic concepts (Bricken et al., 2023; Templeton et al., 2024). However, SAEs alone remain limited; features can still be polysemous, contextually unstable, or overly low-level (Leask et al., 2025; Härle et al., 2025; Peng et al., 2025). Critically, SAEs lack mechanisms for composition and higher-order reasoning.
Formal logic offers explicit compositionality, well-defined inference rules, and transparency in deriving conclusions (Lloyd, 1984; Russell & Norvig, 1995; Shapiro & Kouri Kissel, 2024). Yet, logic-based reasoning presupposes discrete propositional units, whereas LLMs rely on continuous, superposed representations that obscure such structure (Marcus, 2020; Bender et al., 2021; Stammer et al., 2024). While this mismatch has long hindered direct integration, recent progress with SAEs provides a concrete foothold by inducing monosemantic features that approximate discrete propositional units required for symbolic reasoning. To this end, we propose ActivationReasoning (AR), a framework that embeds explicit logical reasoning into the latent space of LLMs. AR operates in three stages (see Fig.  1): (1) Finding latent representations, where atomic concepts are grounded in SAE-derived features; (2) Activating propositions, which monitors their activations during inference; and (3) Logical reasoning, which composes these propositions into higher-order concepts through user-defined rules. This pipeline extends SAEs from feature extraction to structured reasoning and control. As illustrated in Fig.  2, AR can recover composite concepts absent from the SAE space (e.g., Golden Gate Bridge from Bridge, San Francisco, and US); disambiguate polysemous features such as different uses of red; enforce safety constraints directly in latent space; and perform multi-step reasoning. On the analysis side, this enables fine-grained inspection of model behavior, such as tracing failure cases to specific activations, probing compositional generalization, or evaluating adherence to safety rules. On the control side, AR supports direct interventions during inference by amplifying or suppressing activations to steer the model toward desired behaviors.
Our contributions are threefold: (i) we introduce ActivationReasoning (AR), a framework that embeds logical reasoning into the latent space of LLMs by treating SAE features as propositions and composing them through explicit rules; (ii) a new benchmark for reasoning over meta-level concept descriptions, called Rail2Country; and (iii) we demonstrate through experiments on reasoning, meta-concept generalization, and safety that latent activations serve as a viable substrate for structured logical reasoning, enhancing transparency, robustness, and control in LLMs.
@@ 3 ActivationReasoning
ActivationReasoning (AR) enables logical reasoning directly within the latent activations of LLMs (Fig.  1), yielding interpretable, controllable, and robust behavior in downstream tasks. The process has three stages: (i) Finding latent representations, where concept representations are identified in the LLM’s latent space and collected in a dictionary; (ii) Activating propositions, where token-wise activations are detected at inference time and treated as propositional units; and (iii) Logical reasoning, where rules are applied over the activated propositions to infer higher-order structures, compose new concepts, and steer model behavior. As a running example, we infer the concept Golden Gate Bridge, which is not initially present in the dictionary but can be derived from the joint activation of three base concepts: Bridge, San Francisco, and USA.
@@ 3.1 Finding Latent Representations
The first stage of AR (cf. Fig.  1, top half) constructs a set of interpretable concepts grounded in the latent space of the LLM. We first introduce a formal definition of a concept and then explain how the associated representations are constructed.
Formalizing Concepts. We formally define a concept $c$ as a tuple $(n_{c},r_{c},\tau_{c})$. Let $\mathcal{L}\subseteq\mathbb{R}^{d}$ denote the latent space of concept codes produced by a fixed encoder $E$ applied to the model’s hidden states. For each token $t$ with hidden state $h_{t}$, we obtain a latent code via the encoder $\ell_{t}=E(h_{t})\in\mathcal{L}$.
Concretely: (i) $n_{c}$ is a semantic identifier for the concept (e.g., “Bridge”). (ii) $\tau_{c}\in\mathbb{R}_{\geq 0}$ is a soft activation threshold that determines when the concept is considered active. (iii) $r_{c}:\mathcal{L}\to\mathbb{R}_{\geq 0}$ is the concept’s latent representation in this space. Implemented as a function, it maps latent codes $\ell_{t}\in\mathcal{L}$ to a activation score $a(c,t):=r_{c}(\ell_{t})\in\mathbb{R}_{\geq 0}$. The collection of concepts used by AR forms a concept dictionary $\mathcal{D}=\{c_{i}\}_{i=1}^{N}$ containing the building blocks for subsequent logical reasoning.
@@
Toward interpretable activations. Raw LLM activations are highly entangled and subject to superposition, encoding multiple features in overlapping dimensions (Elhage et al., 2022), which obstructs recovery of monosemantic representations (Park et al., 2023). Logical reasoning, however, requires such representations as propositional building blocks. Thus, we project activations into a sparse and more interpretable feature space using SAEs (Bricken et al., 2023), which promote monosemanticity by aligning individual dimensions with distinct, human-interpretable concepts. The framework remains method-agnostic and can directly integrate future advances in interpretability.
@@
Types of latent representations. Building on SAE features, we propose three forms of latent representation $r\in\{\mathcal{R}_{\text{single}},\mathcal{R}_{\text{multi}},\mathcal{R}_{\text{relation}}\}$ (cf. Fig.  1, top right). (1) A Single-feature representation ($\mathcal{R}_{\text{single}}$) ties a concept to one SAE feature, reflecting the common assumption that some features can align closely with human-interpretable notions (Bricken et al., 2023). (2) A Multi-feature representation ($\mathcal{R}_{\text{multi}}$) instead associates a concept with a set of $k$ SAE features. This relaxes the strict monosemantic assumption, acknowledging that features often capture only fragments of a concept or exhibit polysemy (Leask et al., 2025; Härle et al., 2025). Here, features contribute independently through weighted aggregation; no interactions between them are modeled. (3) A Relational-feature representation ($\mathcal{R}_{\text{relation}}$) defines a concept by a set of relations among SAE features. In practice, we instantiate this via a shallow decision tree that applies different thresholds and conditions along different branches, balancing expressiveness with interpretability. Unlike $\mathcal{R}_{\text{multi}}$, this can express structured interactions such as conjunctions (requiring both slurs and stereotypes) or context-dependent exclusions (filtering educational mentions). Simple concepts like Bridge or USA can be captured by $\mathcal{R}_{\text{single}}$ or $\mathcal{R}_{\text{multi}}$, while abstract or contextual notions benefit from $\mathcal{R}_{\text{relation}}$. E.g., the broader concept hate may be recognized if features linked to slurs or demeaning stereotypes are activated, while benign uses of similar words are filtered out. Such patterns require that multiple cues co-occur while others are excluded (e.g., filtering non-toxic educational usage). These structured dependencies are characteristic of $\mathcal{R}_{\text{relation}}$ and cannot be captured by linear aggregation alone.
@@
Concept extraction. Concepts can be obtained either manually or automatically. In the manual case, a human expert defines $r_{c}$ by assigning SAE encoder features $(\mathrm{SAE}_{E})$ to concepts. For example, features consistently activating for the notion of a Bridge can be directly assigned. This direct assignment is used, however, only when a feature is clearly monosemantic, and its link to a concept is unambiguous; in practice, most concepts are constructed via the automatic procedure described next, which scales naturally and does not require manual inspection.
@@ 3.2 Activating Propositions
For $\mathcal{R}_{\text{single}}$, the score corresponds to the value of a single SAE feature (cf. Fig.  1, purple box). For $\mathcal{R}_{\text{multi}}$, evidence is aggregated from multiple indices using a weighting vector $w$ (e.g., log decay), normalized to sum to one, i.e., forming a convex combination. For $\mathcal{R}_{\text{relation}}$, the score is obtained by traversing the decision tree $r_{c}$ on $l_{t}$, producing a probability of concept presence.
@@ 3.3 Logical Reasoning
Reasoning. Having established a set of user-defined rules, reasoning proceeds via forward chaining. Starting from the initial activation matrix $A$ of logical propositions, inference rules are applied iteratively to derive new conclusions, until a fixed point is reached where no further inferences can be drawn (Russell & Norvig, 1995). We formalize the explicit semantics in App.  D. In practice, rules may operate over $A_{\text{local}}$ and $A_{\text{global}}$. The former provides evidence for explicit concepts such as Bridge, San Francisco, and USA at specific tokens, while the latter captures more abstract notions in the sequence, such as sarcasm. For reasoning, we feed discretized activations $A$ into the rule engine (e.g., $A_{\text{local}}[\textit{Bridge},15]=0.9$ is treated as the Boolean proposition Bridge). Applying the compositional rule above to the evidence allows us to infer the higher-level proposition Golden Gate Bridge, even if no SAE feature directly encodes this concept.
Enriched matrix. We denote the resulting enriched matrix of logical propositions as $A^{\prime}$ (cf. Fig.  1, bottom right). In contrast to the raw activations $A$, which contain only directly activated concepts, $A^{\prime}$ integrates both detected and inferred propositions, providing a structured and interpretable representation that can be leveraged for downstream analysis and control.
@@ 3.4 Integration into Downstream Tasks
For analysis, $A^{\prime}$ allows fine-grained inspection of model behavior, such as tracing failure cases to specific concept activations (cf. Sec.  4, Rail2Country), probing compositional generalization (cf. Fig.  2) or evaluating adherence to safety constraints (cf. Sec.  4, safety experiments).
@@ 4 Experimental Evaluations
Experimental Setup. We evaluate AR on two backbone models, Llama-3.1-8B (AI@Meta, 2024) with EleutherAI’s SAE (EleutherAI, 2024) attached after layer 23, and Gemma-2-9B (Team, 2024) with a Gemma-Scope SAE (Lieberum et al., 2024) attached after layer 20. All experiments use greedy decoding. For evaluation, we apply $\mathcal{R}_{\text{multi}}$ in the first three benchmarks and compare $\mathcal{R}_{\text{single}}$, $\mathcal{R}_{\text{multi}}$, and $\mathcal{R}_{\text{relation}}$ in the fourth. We report results against several baselines: the same backbones without AR, larger instruction-tuned models (Llama-3.1-70B and Gemma-2-27B), commercial models (GPT-4o (OpenAI, 2023)), and the reasoning model DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025). Additionally, we provide comparison to further instruction-tuned and chain-of-thought baselines in App.  E and analysis of runtime efficiency in App.  B. Baseline LLMs generate answers directly, whereas AR performs reasoning in the activation space and steers the downstream generation. We provide detailed descriptions of the datasets in App.  A and the hyperparameters in App.  H. We also include an ablation study on the choice of SAE placement, demonstrating that AR remains invariant to this factor (cf. App.  G).
@@
To study this setting, we introduce Rail2Country (R2C), a newly introduced benchmark for reasoning over meta-level concept descriptions, inspired by earlier train-based reasoning tasks (Michalski, 1980; Mitchell, 1997; Helff et al., 2024; 2025). Each instance consists of a context that lists countries with their flag color sequences and a query, a short train description whose sequence of car colors must be mapped to the corresponding country (e.g., red–white–blue $\rightarrow$ France). R2C comes with two versions. In R2C-Mono, colors are explicitly stated in the trains’ descriptions, producing clear SAE activations (Bricken et al., 2023; Templeton et al., 2024). In R2C-Meta, explicit color mentions are replaced with similes (e.g., ’colored like a tomato’ for red). In this way, a color cue is entangled with object semantics (’tomato’), contextual phrasing (’like’), and world knowledge (tomatoes are red). Such descriptions are expected to distribute activation across multiple SAE features (Elhage et al., 2022). Further details and examples are provided in Sec.  A.2.1. For evaluation, we initialize AR with $\mathcal{R}_{\text{multi}}$ using the R2C-Mono train set and encode the context rules into $L$.
@@
Towards Real-World Reasoning Scenarios. Next, we extend our evaluation to ProverQA (Qi et al., 2025), a benchmark of LLM-generated problems that present naturalized reasoning tasks which reflect real-world linguistic diversity. It has three complexity tiers, namely, easy, medium, and hard. Each task provides a context of facts and rules as well as a query that is to be proven true, false, or uncertain. For evaluation, we initialize AR with $\mathcal{R}_{\text{multi}}$ using the set of available propositions, and encode the context rules into $L$. Results on ProverQA are summarized in Tab.  1 (right). We observe that baseline LLMs perform poorly. Larger instruction-tuned models and GPT-4o achieve stronger results on the easy tier but deteriorate on medium and hard problems, and the reasoning model DeepSeek-R1-8B follows a similar trend. In contrast, AR yields substantial gains across all tiers, AR (Llama) reaches 93%/91%/71% (easy/medium/hard) and AR (Gemma) 94%/91%/70%, consistently outperforming both the much larger LLMs and reasoning models. On the hard tier, only AR maintains solid performance near 70%, while all other models fall below 50%. This highlights its robustness to linguistically diverse and complex reasoning tasks and demonstrates that structured reasoning in latent space generalizes effectively to natural, increasingly difficult scenarios.